How many grams of Li2Cr2O7 are in a 10.0 L solution with a molarity of 0.150 M Li2Cr2O7?


1
Expert's answer
2021-02-03T07:07:26-0500

According to the equation:

M = n / V = m / (Mr × V)

where M - molarity, n - number of moles, V - volume, m - mass, Mr - molecular weight

From here:

m = Mr × V × M = 230 g/mol × 10.0 L × 0.150 mol/L = 345 g


Answer: 345 g